# Concurrent-Learning Based Relative Localization in Shape Formation of Robot Swarms

### v1

@@ Appendix B Proof of Theorem 1
$\lVert\tilde{p}_{ij,0}(t_{k})\rVert\leq\left(\sqrt{\left(1-\frac{\lambda_{\min}(S_{ij})^{2}}{(2v_{\max}\Delta t+\lambda_{\max}(S_{ij}))^{2}}\right)}\right)^{k}\lVert\tilde{p}_{ij,0}(t_{0})\rVert,$
which means that $\tilde{p}_{ij,0}$ is global exponentially stable with the convergence rate $a$ defined as (10).
Appendix C Proof of Theorem 2
Define the recorded relative displacement data $u_{ij}(t_{c,m})\triangleq[u_{x,ij}(t_{c,m}),u_{y,ij}(t_{c,m})]^{T}$, as a result, it has
$S_{ij}=\begin{bmatrix}s_{1}&s_{3}\\ s_{3}&s_{2}\end{bmatrix},$ (44)
where $s_{1}$, $s_{2}$ and $s_{3}$ in (44) are defined as $s_{1}=\sum_{m=1}^{\varsigma_{0}ij}u_{x,ij}(t_{c,m})^{2}$, $s_{2}=\sum_{m=1}^{\varsigma_{0}ij}u_{y,ij}(t_{c,m})^{2}$ and $s_{3}=\sum_{m=1}^{\varsigma_{0}ij}u_{x,ij}(t_{c,m})u_{y,ij}(t_{c,m})$. According to Theorem 1 and Lemma 6, the maximum converge rate of the proposed estimator can be achieved if and only if $s_{3}=0$ and $s_{1}=s_{2}\neq 0$. When robot $i$ and robot $j$ perform the designed motion (11), the relative velocity between them is $v_{ij}(t)=[(r_{i}\mathrm{cos}(w_{i}t)-r_{j}\mathrm{cos}(w_{j}t)),(r_{i}\mathrm{sin}(w_{i}t)-r_{j}\mathrm{sin}(w_{j}t))]^{T}$.
Before giving the main proof, we first present several useful equations. With condition (12) is satisfied, one can see that $t_{c,m}=\frac{2\pi}{\varsigma_{0}}(m-1)$. Recall the angular velocity $w_{i}=1/i$ of robot $i$, then it has
$2\sum_{m=1}^{\varsigma_{0}ij}\mathrm{sin}(w_{i}t_{c,m})\mathrm{sin}(w_{j}t_{c,m})$
$=$ $\sum_{m=1}^{\varsigma_{0}ij}\left(\mathrm{cos}(w_{i}t_{c,m}-w_{j}t_{c,m})-\mathrm{cos}(w_{i}t_{c,m}+w_{j}t_{c,m})\right)$
$=$ $\sum_{m=0}^{\varsigma_{0}ij-1}\left(\mathrm{cos}\left(\frac{2\pi(i-j)m}{\varsigma_{0}ij}\right)-\mathrm{cos}\left(\frac{2\pi(i+j)m}{\varsigma_{0}ij}\right)\right).$ (45)
It can be seen that (C) is equal to zero. Similarly, one can check that the following equation also holds when condition (12) is satisfied,
$\sum_{m=1}^{\varsigma_{0}ij}f(w_{i}t_{c,m})f(w_{j}t_{c,m})=0,f\in\{\mathrm{sin},\mathrm{cos}\}$ (46a)
$\sum_{m=1}^{\varsigma_{0}ij}\mathrm{sin}(w_{1}t_{c,m})\mathrm{cos}(w_{2}t_{c,m})=0,w_{1},w_{2}\in\{w_{i},w_{j}\}$ (46b)
$\sum_{m=1}^{\varsigma_{0}ij}\mathrm{sin}(w_{i}t_{c,m})^{2}=\sum_{m=1}^{M}\mathrm{cos}(w_{i}t_{c,m})^{2}\neq 0.$ (46c)
@@ Appendix C Proof of Theorem 2
$u_{x,i}(t_{c,m})$ $=2r_{i}\mathrm{sin}\left(\frac{w_{i}h\Delta t}{2}\right)\mathrm{cos}\left(w_{i}\left(\frac{h\Delta t}{2}+t_{c,m}\right)\right)$
$=s_{i}\mathrm{cos}(w_{i}t_{c,m})-c_{i}\mathrm{sin}(w_{i}t_{c,m}),$ (47a)
$u_{y,i}(t_{c,m})$ $=s_{i}\mathrm{sin}(w_{i}t_{c,m})+c_{i}\mathrm{cos}(w_{i}t_{c,m}),$ (47b)
@@
$s_{3}=$ $\sum_{m=1}^{\varsigma_{0}ij}\left(s_{i}\mathrm{cos}(w_{i}t_{c,m})-c_{i}\mathrm{sin}(w_{i}t_{c,m})-s_{j}\mathrm{cos}(w_{j}t_{c,m})\right.$
$\left.+c_{j}\mathrm{sin}(w_{j}t_{c,m})\right)\left(s_{i}\mathrm{sin}(w_{i}t_{c,m})+c_{i}\mathrm{cos}(w_{i}t_{c,m})\right.$
$\left.-s_{j}\mathrm{sin}(w_{j}t_{c,m})-c_{j}\mathrm{cos}(w_{j}t_{c,m})\right)=0,$
$s_{1}=$ $\sum_{m=1}^{\varsigma_{0}ij}\left(s_{i}\mathrm{cos}(w_{i}t_{c,m})-c_{i}\mathrm{sin}(w_{i}t_{c,m})-s_{j}\mathrm{cos}(w_{j}t_{c,m})\right.$
$\left.+c_{j}\mathrm{sin}(w_{j}t_{c,m})\right)^{2}=\sum_{m=1}^{\varsigma_{0}ij}\left(s_{i}\mathrm{sin}(w_{i}t_{c,m})+c_{i}\mathrm{cos}(w_{i}t_{c,m})\right.$
$\left.-s_{j}\mathrm{sin}(w_{j}t_{c,m})-c_{j}\mathrm{cos}(w_{j}t_{c,m})\right)^{2}=s_{2}\neq 0,$
@@ Appendix D Proof of Theorem 3
According to lemma 1, $\lambda_{\min}(\mathcal{L}+\mathcal{B})>0$. According to (D), the derivative of $V_{l}$ is derived as
$\dot{V}_{l}$ $=\tilde{q}^{T}(\mathcal{L}+\mathcal{B})\dot{\tilde{q}}$
$=-c_{1}\sum_{i=1}^{n}\chi_{i}^{T}\mathrm{sig}\left(\chi_{i}-\delta_{i}\right)^{\alpha},$ (51)
where $\delta{i}\triangleq\sum_{j=1}^{n}a_{ij}\tilde{p}_{ij,0}(t)+\mu_{i}\tilde{p}_{i0,0}(t)\in\mathbb{R}^{2}$. According to Lemma 4, (D) can be simplified as
$\dot{V}_{l}\leq\sum_{i=1}^{n}\left(-\frac{\alpha}{1+\alpha}\lVert\chi_{i}\rVert^{1+\alpha}+\frac{\alpha}{1+\alpha}2^{\frac{1+3\alpha}{2\alpha(\alpha+1)}}\lVert\delta_{i}\rVert^{1+\alpha}\right).$ (52)
@@
$\delta_{i}\leq|\mathcal{N}_{i}|\varepsilon,$
@@
$V_{l}\leq\frac{\frac{\alpha}{1+\alpha}2^{\frac{1+3\alpha}{2\alpha(\alpha+1)}}\left(\sum_{i=1}^{n}(|\mathcal{N}_{i}|\varepsilon)^{1+\alpha}\right)}{(1-\gamma)\frac{\alpha}{1+\alpha}\lambda_{\min}(\mathcal{L}+\mathcal{B})},$ (55)
@@
$\lVert\tilde{q}\rVert\leq b,$

### v4

@@ A-B Proof of Theorem 1
$\lVert\tilde{p}_{ij,0}(t_{k})\rVert\leq\left(\lambda_{ij}\right)^{k}\lVert\tilde{p}_{ij,0}(t_{0})\rVert$
where $\lambda_{ij}$ is defined in (9)
A-C Proof of Theorem 2
Define the recorded relative displacement data $u_{ij}(t_{{\rm c},m})\triangleq[u_{x,ij}(t_{{\rm c},m}),u_{y,ij}(t_{{\rm c},m})]^{T}$, as a result, the data matrix $S_{ij}$ is $S_{ij}=\begin{bmatrix}s_{1}&s_{3}\\ s_{3}&s_{2}\end{bmatrix}$, where $s_{1}$, $s_{2}$ and $s_{3}$ are defined as $s_{1}=\sum_{m=1}^{\varsigma_{0}ij}u_{x,ij}(t_{{\rm c},m})^{2}$, $s_{2}=\sum_{m=1}^{\varsigma_{0}ij}u_{y,ij}(t_{{\rm c},m})^{2}$ and $s_{3}=\sum_{m=1}^{\varsigma_{0}ij}u_{x,ij}(t_{{\rm c},m})u_{y,ij}(t_{{\rm c},m})$. According to Theorem 1 and Lemma 6, the maximum converge rate of the proposed estimator can be achieved if and only if $s_{3}=0$ and $s_{1}=s_{2}\neq 0$. When robot $i$ and robot $j$ perform the designed motion (10), the relative velocity between them is $v_{ij}(t)=[(r_{i}\mathrm{cos}(w_{i}t)-r_{j}\mathrm{cos}(w_{j}t)),(r_{i}\mathrm{sin}(w_{i}t)-r_{j}\mathrm{sin}(w_{j}t))]^{T}$. Before giving the main proof, we first present several useful equations. With condition (11) is satisfied, one can see that $t_{{\rm c},m}=\frac{2\pi}{\varsigma_{0}}(m-1)$. Recall the angular velocity $w_{i}=1/i$ of robot $i$, then it has
$2\sum_{m=1}^{\varsigma_{0}ij}\mathrm{sin}(w_{i}t_{{\rm c},m})\mathrm{sin}(w_{j}t_{{\rm c},m})$ (36)
$=$ $\sum_{m=1}^{\varsigma_{0}ij}\left(\mathrm{cos}(w_{i}t_{{\rm c},m}-w_{j}t_{{\rm c},m})-\mathrm{cos}(w_{i}t_{{\rm c},m}+w_{j}t_{{\rm c},m})\right)$
$=$ $\sum_{m=0}^{\varsigma_{0}ij-1}\left(\mathrm{cos}\left(\frac{2\pi(i-j)m}{\varsigma_{0}ij}\right)-\mathrm{cos}\left(\frac{2\pi(i+j)m}{\varsigma_{0}ij}\right)\right).$
It can be seen that (36) is equal to zero. One can check that the following equation holds when condition (11) is satisfied,
$\sum_{m=1}^{\varsigma_{0}ij}f(w_{i}t_{{\rm c},m})f(w_{j}t_{{\rm c},m})=0,f\in\{\mathrm{sin},\mathrm{cos}\}$ (37a)
$\sum_{m=1}^{\varsigma_{0}ij}\mathrm{sin}(w_{1}t_{{\rm c},m})\mathrm{cos}(w_{2}t_{{\rm c},m})=0,w_{1},w_{2}\in\{w_{i},w_{j}\}$ (37b)
$\sum_{m=1}^{\varsigma_{0}ij}\mathrm{sin}(w_{i}t_{{\rm c},m})^{2}=\sum_{m=1}^{M}\mathrm{cos}(w_{i}t_{{\rm c},m})^{2}\neq 0.$ (37c)
@@ A-C Proof of Theorem 2
$u_{x,i}(t_{{\rm c},m})$ $=2r_{i}\mathrm{sin}\left(\frac{w_{i}h\Delta t}{2}\right)\mathrm{cos}\left(w_{i}\left(\frac{h\Delta t}{2}+t_{{\rm c},m}\right)\right)$
$=s_{i}\mathrm{cos}(w_{i}t_{{\rm c},m})-c_{i}\mathrm{sin}(w_{i}t_{{\rm c},m})$ (38a)
$u_{y,i}(t_{{\rm c},m})$ $=s_{i}\mathrm{sin}(w_{i}t_{{\rm c},m})+c_{i}\mathrm{cos}(w_{i}t_{{\rm c},m})$ (38b)
@@
$s_{3}=$ $\sum_{m=1}^{\varsigma_{0}ij}\left(s_{i}\mathrm{cos}(w_{i}t_{{\rm c},m})-c_{i}\mathrm{sin}(w_{i}t_{{\rm c},m})-s_{j}\mathrm{cos}(w_{j}t_{{\rm c},m})\right.$
$\left.+c_{j}\mathrm{sin}(w_{j}t_{{\rm c},m})\right)\left(s_{i}\mathrm{sin}(w_{i}t_{{\rm c},m})+c_{i}\mathrm{cos}(w_{i}t_{{\rm c},m})\right.$
$\left.-s_{j}\mathrm{sin}(w_{j}t_{{\rm c},m})-c_{j}\mathrm{cos}(w_{j}t_{{\rm c},m})\right)=0$
$s_{1}=$ $\sum_{m=1}^{\varsigma_{0}ij}\left(s_{i}\mathrm{cos}(w_{i}t_{{\rm c},m})-c_{i}\mathrm{sin}(w_{i}t_{{\rm c},m})-s_{j}\mathrm{cos}(w_{j}t_{{\rm c},m})\right.$
$\left.+c_{j}\mathrm{sin}(w_{j}t_{{\rm c},m})\right)^{2}=\sum_{m=1}^{\varsigma_{0}ij}\left(s_{i}\mathrm{sin}(w_{i}t_{{\rm c},m})+c_{i}\mathrm{cos}(w_{i}t_{{\rm c},m})\right.$
$\left.-s_{j}\mathrm{sin}(w_{j}t_{{\rm c},m})-c_{j}\mathrm{cos}(w_{j}t_{{\rm c},m})\right)^{2}=s_{2}\neq 0$
@@ A-D Proof of Theorem 3
According to lemma 1, $\lambda_{\min}(\mathcal{L}+\mathcal{B})>0$. According to (A-D), the derivative of $V_{l}$ is derived as
$\dot{V}_{l}=\tilde{q}^{T}(\mathcal{L}+\mathcal{B})\dot{\tilde{q}}=-c_{1}\sum_{i=1}^{n}\chi_{i}^{T}\mathrm{sig}\left(\chi_{i}-\delta_{i}\right)^{\alpha}$ (42)
where $\delta{i}\triangleq\sum_{j=1}^{n}a_{ij}\tilde{p}_{ij,0}(t)+\mu_{i}\tilde{p}_{i0,0}(t)\in\mathbb{R}^{2}$. According to Lemma 4, (42) can be simplified as
$\dot{V}_{l}\leq\sum_{i=1}^{n}\left(-\frac{\alpha}{1+\alpha}\lVert\chi_{i}\rVert^{1+\alpha}+\frac{\alpha}{1+\alpha}2^{\frac{1+3\alpha}{2\alpha(\alpha+1)}}\lVert\delta_{i}\rVert^{1+\alpha}\right).$
@@
$\delta_{i}\leq|\mathcal{N}_{i}|\varepsilon$
@@
$V_{l}\leq\frac{\frac{\alpha}{1+\alpha}2^{\frac{1+3\alpha}{2\alpha(\alpha+1)}}\left(\sum_{i=1}^{n}(|\mathcal{N}_{i}|\varepsilon)^{1+\alpha}\right)}{(1-\gamma)\frac{\alpha}{1+\alpha}\lambda_{\min}(\mathcal{L}+\mathcal{B})}$ (45)
@@
$\lVert\tilde{q}\rVert\leq b$
@@
Appendix B Technical Details
B-A Additional simulation results
Figure 9: A. Trajectory and final position of the swarm with 25, 50, 100 and 200 robots. B. Coverage rate, entering rate and distribution uniformity of the robot swarms to form the mathematical sum symbol.
Figure 10: A. Trajectory and final position of the swarm with 25, 50, 100 and 200 robots. B. Coverage rate, entering rate and distribution uniformity of the robot swarms to form the methane molecule shape.
[…]
B-B Discussion on parameter setting
TABLE II: Parameter list
Simulation & Experiment $\Delta t$(s) $h$ $\lambda_{0}$ $c_{1}$ $\alpha$ $\delta_{t}$(s) $\delta_{0}$(m) $n_{\mathrm{cell}}$ $\kappa_{1}$ $\kappa_{2}$ $\kappa_{3}$ $\kappa_{4}$
Simulation: dark shape 0.05 40 0.1 0.1 0.5 1 0.01 2529 10 15 25 2
Simulation: letter “R” shape 0.05 40 0.1 0.1 0.5 1 0.01 2470 10 15 25 2
Simulation: “sum” symbol shape 0.05 40 0.1 0.1 0.5 1 0.01 2130 10 15 25 2
Simulation: “molecule” shape 0.05 40 0.1 0.1 0.5 1 0.01 2251 10 15 25 2
Experiment: “arrow” shape 0.05 40 0.2 0.1 0.5 1 0.02 2209 10 15 23 2
Experiment: letter “T” shape 0.05 40 0.2 0.1 0.5 1 0.02 2266 10 15 23 2
We explain in detail all the parameters involved in the method proposed in this article and their corresponding meanings below. In addition, for all our simulation and hardware experimental scenarios, we have discussed the parameter selection method and the sensitivity of the proposed method to the parameters. The main parameters adopted in our methods can be classified into two classes, one class is the relative localization and shape agreement related, including $\Delta t$, $h$, $\lambda_{0}$, $c_{1}$, $\alpha$, $\delta_{t}$ and $\delta_{0}$. The other class is related to the shape formation control, including $n_{\mathrm{cell}}$, $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$ and $\kappa_{4}$.
For the parameters adopted in the relative localization and shape localization agreement:
  • The parameter $\Delta t$ represents the measurement time interval of the sensor, and $h$ represents the ratio of the time interval for collecting data to the measurement time interval of the sensor. The selections of $\Delta t$ and $h$ are generally related to the robot’s motion speed and the IO measurement noise. In order to prevent the robot’s motion displacement per unit time ($\Delta t$ multiplied by $h$) from being submerged in the measurement noise of the IO, we recommend that the norm of the robot’s displacement per unit time be more than $10$ times the average measurement noise of the odometer. Below we provide a calculation example. Taking the Tb3 robot onboard odometer used in this work as an example, the odometer error (norm of the displacement measurement error) is generally less than $0.01$m within a short period of time (10 seconds) through our extensive practical verification. For the robot making localization-enhancement motion, its velocity is mostly within the range of 0.05 and 0.22m/s (maximum velocity of Tb3 robot). To ensure that the displacement of the robot per unit time ($\Delta t$ multiplied by $h$) is more than $10$ times the average noise level, the product of $\Delta t$ and $h$ is recommended to be around $2$. As a result, the norm of the robot’s displacement per unit time ($\Delta t$ multiplied by $h$) can be calculated roughly within the range of $0.1$ to $0.44$.
  • The parameter $\lambda_{0}$ denotes the threshold for the ratio $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ of the collected data matrix $S_{ij}$. When $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ is greater than $\lambda_{0}$, the localization-enhancement motion stop. $\lambda_{0}$ measures the amount of information collected, the larger $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$, the faster the relative position estimator converges. For circular motion (radius between $0.3$m and $1$m) of robots at different angular velocities (between $0.1$m/s and $1$m/s), typically $\frac{\lambda_{\min}(S_{ij})}{\lambda_{\max}(S_{ij})}$ can reach around $0.2$. As a result, $\lambda_{0}$ is recommended to be around $0.1$ in practical applications to achieve faster convergence speed.
  • The parameters $c_{1}$ and $\alpha$ are the gain and exponent of the seed robot relative position estimator, respectively. For $c_{1}$, generally it can be selected as a constant greater than 00 but less than $10$. A large number of consensus related works are also selected in this way. In our method, $\alpha$ is a parameter within the range of 00 and $1$, as similar to most finite-time consensus results. In our method, it is shown that the shape agreement error $\tilde{q}$ will converge into an adjustable set $\Omega=\{\lVert\tilde{q}\rVert\leq b\}$ in a finite time $t_{a}+t_{l}$ with the proposed estimation law, where $b$, $t_{a}$ and $t_{l}$ are defined as
    $b=\sqrt{\frac{2^{\frac{1+3\alpha}{2\alpha(\alpha+1)}}\left(\sum_{i=1}^{n}(|\mathcal{N}_{i}(t_{0})|\varepsilon)^{1+\alpha}\right)}{(1-\gamma)\lambda_{\min}(\mathcal{L}+\mathcal{B})^{2}}}$ (46)
    $t_{l}=\frac{2(1+\alpha)V_{l}(t_{a})^{\frac{1-\alpha}{2}}}{\lambda_{\min}(\mathcal{L}+\mathcal{B})\gamma(1-\alpha)}$ (47)
    $t_{a}=\mathrm{\max}\{\frac{\mathrm{ln}(\varepsilon\lVert\tilde{p}_{ij,0}(t_{0})\rVert^{-1})}{\mathrm{ln}(\lambda_{ij})}\Delta t,(i,j)\in\mathcal{G}(t_{0})\}$ (48)
    As far as the convergence time $t_{l}$ is concerned, it can be seen that the smaller $\alpha$, the smaller the convergence time $t_{l}$. However, the smaller $\alpha$, the larger $b$, which means that the agreement error $\tilde{q}$ converges to a larger range in a faster time. So $\alpha$ should not be too small (approaching 0) nor too large (approaching 1). Through our extensive practical verification, $\alpha$ is recommended to be around $0.5$. According to Eq. (18) in the revised manuscript, the smaller the value of $\alpha$, the faster the convergence speed of the estimator.
  • The parameters $\delta_{t}$ and $\delta_{0}$ denote the unit time of the seed robot relative position estimator and threshold of estimator change rate, respectively. For the seed robot relative position estimator, the rate $\frac{\lVert\hat{q}_{i,0}(t)-\hat{q}_{i,0}(t-\delta_{t})\rVert}{\delta_{t}}$ reflects the changes in estimator values per unit time $\delta_{t}$. When $\frac{\lVert\hat{q}_{i,0}(t)-\hat{q}_{i,0}(t-\delta_{t})\rVert}{\delta_{t}}$ is greater than $\delta_{0}$, the agreement process can be considered as over. Generally, $\delta_{t0}$ can be easily chosen as $1$s, while for general precise requirements, $\delta_{0}$ can be chosen at around $0.01$, indicating that the estimator change rate is less than $1$ centimeters per second.
For the parameters adopted in the distributed shape formation control method:
  • The parameter $n_{\mathrm{cell}}$ denotes the black cell number in the user-given binary image. The larger the value of $n_{\mathrm{cell}}$, the clearer the target image. Generally, $40\times 50$ image can be used to represent most complex shapes, so $n_{\mathrm{cell}}$ is recommended to be around $2000$.
  • The parameters $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$ and $\kappa_{4}$ denote the shape-entering command gain, shape-exploration command gain, collision-avoidance gain and velocity-alignment gain, respectively. The larger $\kappa_{1}$, the greater the force that pushes the robot into shape. The larger $\kappa_{2}$, the greater the force exerted by the robot in shape exploration. The larger $\kappa_{3}$, the greater the collision avoidance force between robots. The larger $\kappa_{4}$, the more obvious the speed alignment effect between robots. Through our extensive practical verification, $\kappa_{1}$, $\kappa_{2}$ and $\kappa_{3}$ are chosen at around $10$, and $\kappa_{4}$ is recommended to be around $5$.
The detailed parameters in our simulation and hand-ware experiments can be found in Table II. From the parameter value in Table II, it can be seen that our estimation algorithm and control algorithm do not make significant adjustments to parameters in simulations and hardware experiments with different target shapes and robot swarm scales. In other words, the algorithm proposed in this paper has low sensitivity to these parameters.